MARR=10% Alt A Alt B
Initial Cost ($1,500) ($1,200)
Annual Rev $1,200 $1,500
Annual Cost ($250) ($400)
Salvage Val $400 $ 100.00
Useful Life 12 9
AW

find the annual worth of project A and B. Use repeatability assumption
Answer #1

A

B

Initial Investment

​($1,500)

(​$1,200)

Annual Revenue

​$1,200

​$1,500

Annual Cost

​($250)

​($400)

Salvage Value

​$400

​$100

Useful Life

12 years

9 years

MARR = 10%

Find the AW of both the alternatives.

Both the alternatives have unequal lives. We have to convert the unequal lives into equal life by using the repeatability assumption. Calculate the LCM of 12 years of A and 9 years of B. The LCM is 36 years. The common life will be 36 years for both the alternatives.

Therefore, the alternative A is to be repeated 3 times to make it equal to 36 years.

Similarly, the alternative B is to be repeated 4 times to make it equal to 36 years.

Since the life is not equal, we need to calculate the PW and then we can calculate the Annual Worth.

Initial Cost = 1500

Annual Revenue = 1200

Annual cost = 250

Net Annual Revenue = 1200 – 250 = 950

Salvage Value = 400

Step 1 Calculate PW

PW = -1500 – 1500 (P/F, 10%, 12) – 1500 (P/F, 10%, 24) + 950 (P/A, 10%, 36) + 400 (P/F, 10%, 12) + 400 (P/F, 10%, 24) + 400 (P/F, 10%, 36)

PW = -1500 – 1500 (0.31863) – 1500 (0.10153) + 950 (9.67651) + 400 (0.31863) + 400 (0.10153) + 400 (0.03235) = 7243

Step 2 Calculate AW

AW = PW (A/P, 10%, 36)

AW = 7243 (0.10334) = 748.5 or 749

Alternative B

Initial Cost = 1200

Annual Revenue = 1500

Annual cost = 400

Net Annual Revenue = 1500 – 400 = 1100

Salvage Value = 100

Step 1 Calculate PW

PW = -1200 – 1200 (P/F, 10%, 9) – 1200 (P/F, 10%, 18) – 1200 (P/F, 10%, 27) + 1100 (P/A, 10%, 36) + 100 (P/F, 10%, 9) + 100 (P/F, 10%, 18) + 100 (P/F, 10%, 27) + 100 (P/F, 10%, 36)

PW = -1200 – 1200 (0.42410) – 1200 (0.17986) – 1200 (0.07627) + 1100 (9.67651) + 100 (0.42410) + 100 (0.17986) + 100 (0.07627)   + 100 (0.03235) = 8699

Step 2 Calculate AW

AW = PW (A/P, 10%, 36)

AW = 8699 (0.10334) = 899

Annual Worth of Alternative A 743

Annual Worth of Alternative B 899